y'''+3y''+3y'+y=0


1
Expert's answer
2021-05-03T06:59:24-0400

Given the differential equation: "y'''+3y''+3y'+y=0"

Require to find the general solution of the given differential equation.

An Auxiliary Equation corresponding to the given homogeneous linear differential equation is

"m^3+3m^2+3m+1=0"

"\Rightarrow (m+1)^3=0"

"\Rightarrow m=-1,-1,-1"

The general solution to the given differential equation is

"y(x)=(c_1+c_2x+c_3x^2)e^{-x}"

Therefore, general solution to the given differential equation is

"y(x)=(c_1+c_2x+c_3x^2)e^{-x}"